OpenStudy (anonymous):

use the discriminant to find the number of real solutions of the following quadratic equation: 2x^2+2x+2=0

OpenStudy (aum):

First step: Divide throughout by 2 to simplify.

OpenStudy (anonymous):

just put your a ,b and c in the form\[\pm\sqrt{b^{2}-4ac}\] and then use the table in the link http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php

OpenStudy (anonymous):

x^2+x+1=0

OpenStudy (aum):

Yes. Compare x^2+x+1=0 to ax^2 + bx + c = 0 and identify a, b and c. Then find the discriminant b^2 - 4ac. If the discriminant < 0, no real solutions If the discriminant = 0, one real solution (twice) If the discriminant > 0, two real solutions